Can anyone answer the following equation please?
9 = -3 + n + 2n

Answer by Alan3354(69443)

Can anyone answer the following equation please?
9 = -3 + n + 2n
-----------------
The equation is not calling on the phone, or asking a question.
----------
Try this -
Solve for n:
9 = -3 + n + 2n
9 = -3 + 3n
Add 3
12 = 3n
Divide by 3
n = 4